### v1

# AETAS: Analysis of Evolving Temporal Affect and Semantics for Legal History

Abstract
Digital-humanities work on semantic shift often alternates between handcrafted close readings and opaque embedding machinery. We present a reproducible expert-system style pipeline that quantifies and visualises lexical drift in the Old Bailey Corpus (1720–1913), coupling interpretable trajectories with legally meaningful axes. We bin proceedings by decade with dynamic merging for low-resource slices, train skip-gram embeddings, align spaces through orthogonal Procrustes, and measure both geometric displacement and neighborhood turnover. Three visual analytics outputs, which are drift magnitudes, semantic trajectories, and movement along a mercy-versus-retribution axis, expose how justice, crime, poverty, and insanity evolve with penal reforms, transportation debates, and Victorian moral politics. The pipeline is implemented as auditable scripts so results can be reproduced in other historical corpora.
keywords:
Semantic drift , Diachronic embeddings , Digital humanities , Old Bailey , Expert systems[inst1]organization=PingCAP, Data & AI-Innovation Lab,city=Beijing,country=China
Highlights
Expert-system pipeline: decadal binning with adaptive merging, SGNS training, Procrustes alignment, and scripted visual analytics.
Digital-humanities storyline: connects lexical drift to penal transportation, the rise of professionalised justice, and Victorian moral care.
New interpretability device: a mercy-retribution axis quantifies value-laden shifts in legal vocabulary.
Reproducibility: all steps (binning, training, alignment, drift scoring, visualisation) are scripted for reruns on new corpora.
1 Introduction
Semantic drift studies in digital humanities now leverage distributional semantics to explain cultural change [1, 2]. Yet domain experts often face two gaps: (1) storylines that connect numeric drift to socio-legal events, and (2) tooling that behaves like an expert system with auditable steps and repeatable outputs. We address these gaps through AETAS (Analysis of Evolving Temporal Affect and Semantics; Latin aetas, meaning era or lifetime), a focused case study on the Old Bailey Corpus (OBC), a canonical source for London courtroom proceedings (1720–1913). The goal is to surface how legal and social concepts moved from early modern courts to late-Victorian professional justice.
We contribute: (i) a reproducible pipeline that treats diachronic embeddings as chained inference steps, (ii) a narrative that aligns semantic displacement with legal history, and (iii) visual artefacts tuned for interpretability rather than black-box accuracy.
2 Related work
Early semantic-change models used frequency- and context-based measures [1]. Subsequent surveys emphasise alignment stability, evaluation difficulties, and downstream interpretability [2, 3]. Contextualised models and probing have been explored for finer-grained senses [4]. Digital-humanities applications couple embeddings with historical argumentation, e.g., criminality, gender, and class [8, 11]. Our work extends this line with an expert-system stance: each step (binning, alignment, drift scoring) is explicit, audited, and paired with narrative-ready visualisations.
3 Data and methods
3.1 Corpus and temporal binning
We use OBC 2.0 TEI-XML, converting proceedings to a parquet corpus. Text is lowercased, punctuation stripped except hyphens, and grouped into 10-year bins. Bins with fewer than five million tokens merge forward (resulting in an initial 1674s bin through 1749), preserving chronological order.
3.2 Embedding training
Skip-gram with negative sampling (vector size 300, window 5, min count 10, negative samples 10, epochs 5, seed 42) is trained per bin using gensim [12]. Tokens per bin form sentence streams, avoiding shuffling across periods.
3.3 Space alignment
To compare meanings, we align each target space to a 1900s base using orthogonal Procrustes on the shared vocabulary [1]. Let and be base and target matrices; the rotation is obtained from with , ensuring distances are preserved.
3.4 Drift scoring and neighbors
Semantic drift between a word vector at start and at end is . We also track neighborhood turnover via Jaccard overlap between top- neighbors at the two times.
3.5 Normative axis
To ground interpretability, we build a mercy-retribution axis using averaged vectors of {mercy, pity, charity, kindness} minus {punishment, law, authority, order, duty}. Projections of target terms on this axis capture value-laden movement beyond raw drift.
3.6 Reproducibility
All stages (corpus parsing, binning, SGNS training, alignment, drift scoring, figure generation) are scripted for repeatable execution on other corpora; code and configuration are available in the project repository.
3.7 Corpus diagnostics
Table 1 reports decadal statistics. Early sparse decades (1674–1749) merge to satisfy the 5M-token rule; peak coverage is 1840s (13.2M tokens, 85K types). A naive lemmatisation heuristic reduces vocabulary size by 11–15% across decades, indicating moderate orthographic fragmentation without dominating the counts.
| Bin | Years | Tokens (M) | Vocab (K) |
| 1674s | 1674–1749 | 5.72 | 56.6 |
| 1750s | 1750–1769 | 5.16 | 40.6 |
| 1770s | 1770–1789 | 9.51 | 51.1 |
| 1790s | 1790–1799 | 6.09 | 42.3 |
| 1800s | 1800–1809 | 5.67 | 40.5 |
| 1810s | 1810–1819 | 5.60 | 46.6 |
| 1820s | 1820–1829 | 8.40 | 54.4 |
| 1830s | 1830–1839 | 12.38 | 76.9 |
| 1840s | 1840–1849 | 13.18 | 85.1 |
| 1850s | 1850–1859 | 9.99 | 59.0 |
| 1860s | 1860–1869 | 8.79 | 59.5 |
| 1870s | 1870–1879 | 7.98 | 58.5 |
| 1880s | 1880–1889 | 9.57 | 60.4 |
| 1890s | 1890–1899 | 7.46 | 55.1 |
| 1900s | 1900–1909 | 8.41 | 57.6 |
| 1910s | 1910–1913 | 2.59 | 32.5 |
4 Results
4.1 Drift magnitudes
Figure 1 shows cosine drift per target word across spans. On average, drift is , with the largest shifts in justice (0.57 in 1750s–1850s) and transportation (0.53), reflecting the decline of deportation and the rise of professional judges. Social terms such as charity exhibit long-horizon drift (0.50) as philanthropy professionalises.
Table 2 summarises the largest displacements and sparse neighbor overlap, highlighting that shifts are driven by contextual replacement rather than stable colexification.
| Word | Span | Domain | Drift | Overlap |
| justice | 1750s–1850s | legal | 0.565 | 0.000 |
| transportation | 1750s–1850s | social | 0.533 | 0.000 |
| justice | 1750s–1900s | legal | 0.520 | 0.000 |
| charity | 1750s–1900s | social | 0.500 | 0.000 |
| insanity | 1750s–1850s | social | 0.460 | 0.000 |
| insanity | 1750s–1900s | social | 0.427 | 0.000 |
| charity | 1850s–1900s | mix | 0.401 | 0.000 |
| crime | 1750s–1900s | legal | 0.388 | 0.111 |
4.2 Semantic trajectories
Figure 2 plots PCA projections of aligned vectors. Justice moves from a dispersed early-modern sense (neighbors include baron, recorder) toward a tight legal-professional cluster (bigham, phillimore). Poverty shifts away from moral failing toward structural conditions (neighbors like ill-health, schemes in 1900s).
4.3 Value-laden movement
Figure 3 shows projections on the mercy-retribution axis. Justice drifts slightly toward a more punitive neighbourhood by 1900s (score ), consistent with late-Victorian emphasis on deterrence despite procedural reforms. Insanity and poverty trend toward the merciful side by 1850s before softening in 1900s, echoing medicalisation and social-reform rhetoric. Figure 5 adds seed-variation bands; the sign of justice is robust to leave-one-out seeds, while poverty shows wider sensitivity because it co-occurs with both moralising and structural contexts.
4.4 Robustness checks
To mitigate brittleness, we enlarged neighborhoods to (Figure 4). Jaccard overlap rises modestly (mean 0.044 vs. 0 for ), indicating that low overlap is not solely a small- artifact. Target-word frequencies per million tokens stay within the same order across bins (e.g., justice 68290), suggesting drift is not driven by single bursts. A naive lemmatisation lowers type counts by 11–15%, implying that spelling variation inflates vocabulary moderately but does not dominate.
4.5 Contextual, CCA, VecMap, and pivot baselines
We ran a contextual BERT-base baseline (up to 120 contexts/bin; early bins are sparse) with APD over contextual embeddings and substitute-based scaled JSD. Transportation shows the largest APD and JSD in 1750s–1850s, mirroring the SGNS signal; justice APD is moderate. A CCA alignment baseline (50 components) yields drift for justice and transportation that aligns with SGNS trends, with mean canonical correlations 0.63–0.68 over 6k–10k shared types. VecMap unsupervised alignment (capped vocabulary) produces consistent rankings (e.g., justice high drift in 1750s–1850s). To avoid explicit alignment, we compute a pivot baseline: rank similarity to 500 high-frequency pivots per bin and compare top-50 lists; pivot Jaccard remains low for justice but higher for crime/poverty, with mid-level Spearman stability. A temporal-reference norm baseline (anchor 1900s) also shows large norms for justice and transportation, consistent with long-range drift.
4.6 Incremental SGNS and multi-seed variance
We trained an incremental (time-forward) SGNS with three seeds over 1750s1850s1900s (vector size 200, 3 epochs per bin), saving snapshots per bin. Drift means (std across seeds) are: justice 0.40 (0.03) for 1750s–1850s, 0.50 (0.02) for 1750s–1900s; transportation 0.14 (0.01) for 1750s–1850s; crime 0.30 (0.01) for 1750s–1900s. Variance is low (std ) indicating seed robustness.
4.7 Split-half stability
Splitting each bin (1750s, 1850s, 1900s) into halves, training SGNS on each half (20 repeats), and measuring within-bin drift shows substantial sampling variance: justice mean split-half drift 0.71 (std up to 0.10), crime 0.63 (std 0.08), transportation 0.72 (std 0.08) (Figure 6(a)). We therefore report a net drift (Figure 7) that subtracts the average split-half baseline and a standardized effect size (net divided by pooled split-half std); only charity (1750s–1900s) remains positive, with a small .
4.8 Frequency controls and preprocessing
Regressing SGNS drift on mean and delta frequency across spans yields no significant coefficients (p0.09, ); contextual APD and incremental SGNS show similarly non-significant frequency effects (p0.14, ), suggesting observed drift is not explained by frequency trends alone. A lightweight NER/POS filter on sampled texts removes 10–16% of tokens and 21–35% of types (Figure 6(b)), indicating named entities and numerals contribute noticeable noise. Preprocessing: lowercase, strip punctuation except hyphens, no stopword removal, no lemmatization in the main runs.
4.9 NER/POS filtering sanity check
Applying a lightweight NER/POS filter (spaCy en_core_web_sm) to 200-sample texts per bin removes 10–16% of tokens and 21–35% of types. This suggests that proper-noun/number noise is non-trivial; integrating NER/POS-aware preprocessing is left for future modeling runs.
5 Discussion
The results align qualitatively with digital-humanities narratives: (1) erosion of transportation in British penal practice; (2) professionalisation of judges and a move from status-based justice to codified procedure; (3) medicalisation of madness and insanity. Yet split-half instability is large, and net drifts after subtracting that baseline are mostly small or negative (Figure 7), so these findings should be read as hypothesis-generating rather than definitive evidence of semantic change.
Robustness additions partially address instability: expanding reduces but does not eliminate neighborhood sparsity (Figure 4); seed-variation bands suggest the mercy-retribution sign of justice is robust (Figure 5); corpus and frequency diagnostics guard against coverage/frequency artifacts; net-drift/z heatmaps contextualize effect sizes (Figure 7). Remaining risks include anchoring bias from using 1900s as the sole base, limited contextual baselines in sparse bins, uncalibrated value axes (no external labels), and proper-noun noise.
6 Limitations and future work
Open items:
Temporal chaining/reference alignments beyond the 1900s anchor, plus OT/hubness mitigation and dynamic smoothing.
Bootstrapped stability for drift and axes; hubness diagnostics; token-based contextual baselines with balanced sampling/PRT.
Calibrated value axes with external labels (e.g., sentencing outcomes or expert ratings), additional moral/procedural axes, richer seed bootstraps, and integrating NER/POS filtering into main runs.
Stronger polysemy/context-diversity controls following [7].
Evaluating semantic change remains challenging without gold standards [3]; crowdsourced historical judgments or links to sentencing outcomes would strengthen validation and the “expert system” claims.
7 Conclusion
We demonstrated an auditable semantic-drift pipeline that couples diachronic embeddings with digital-humanities storytelling. Each module is an inference component whose outputs can be inspected and recombined. Given the high within-bin variance and modest net effects, the current claims should be viewed as hypotheses; with stronger calibration (chain alignments, contextual baselines, calibrated value axes, and human/external validation), the approach can mature into a reusable template for domain-specific semantic change analysis.